ordered: Aminophylline IV infusion for a patient with acute asthma. Titrate 0.3 to 0.6 mg/kg/hr. your patient weighs 75kg. The drug comes premixed as Aminophylline 500mg/250mL. what ml/hr flow rate will be set on the infusion pump for 0.3mg/kg/hr?
Answer #1

ANSWER.

* Calculate the flow rate(ml/hr) will be set on the infusion pump for 0.3mg/kg/hr.

          Formula

               Flow rate = Total volume(ml) /Time(hr).

       Calculation of total volume(ml).

             Formula = Desired dose(D) /Dose in hand(H) * Quantity(Q).

       Here

           Ordered dose = 0.3 mg/kg/hr.

     Weight of the patient = 75 kg.

             Desired dose(D) for a patient weight 75 kg = 0.3 * 75

                                                                                  = 22.5 mg/hr.

                     Dose in hand (H) = 500 mg.

                        Quantity(Q) = 250 ml.

    As per above formula

            Total volume(ml) of medication required = D/H * Q

                                                                             = 22.5 / 500 * 250

                                                                             = 0.045 * 250

                                                                             = 11.25 ml.

                                                   Time (hr) = 1 hour.

       So

           Flow rate(ml/hr) set on infusion pump = 11.25 ml/hr.